Two cubes each with 6 cm edge are joined end to end. The surface area of the resulting cuboid is ____

Grade:12th pass

1 Answers

Pawan Prajapati
2 years ago
Answer: Explanation: [360 cm²] Hint: Length of resulting cuboid = 12 cm. Width = 6 cm, height = 6 cm. ∴ Surface area = 2(12 × 6 +12 × 6 + 6 × 6) = 2(180) = 360 cm².